OpenStudy (anonymous):

Algebra 2 please help, I'll give you a medal :)

OpenStudy (anonymous):

Where's the question lol

OpenStudy (anonymous):

Given a geometric sequence in the table below, create the explicit formula and list any restrictions to the domain. n a_n 1 -4 2 20 3 -100 A. a_n = −5(−4)^(n − 1) where n ≥ 1 B. a_n = −4(−5)^(n − 1) where n ≥ 1 C. a_n = −4(5)^(n − 1) where n ≥ −4 D. a_n = 5(−4)^(n − 1) where n ≥ −4

hartnn (hartnn):

you know now how to find common ratio ?

OpenStudy (anonymous):

No?

hartnn (hartnn):

common ratio r = 2nd term /1st term = 3rd term /2nd term = ...and so on now can you find ?

OpenStudy (anonymous):

I'm not sure but I think the answer is A..

hartnn (hartnn):

its not A

hartnn (hartnn):

the general term is \(\large a_n = a_1 r^{n-1}\) a1 = 1st term r= common ratio